OpenStudy (anonymous):

the product of two numbers is 77, the sum of the two numbers is 18. find the two numbers.

OpenStudy (anonymous):

11 and 7

OpenStudy (anonymous):

s11and 7

OpenStudy (anonymous):

x * y = 77 x + y = 18 use this idea

OpenStudy (anonymous):

x*y = 77 y = (77/x) x+y = 18 x+(77/x) = 18 x^2 - 18x + 77 = 0 solve the above x = 11,7 So, 2 numbers are 11 & 7